OpenStudy (lilmike234):

The vertex of a parabola is (3,2). A second point on the parabola is (1,5). Which of the following points is also on the graph? A. (-1,5) B. (3,-2) C. (3,7) D. (5,5)

OpenStudy (mertsj):

To get from the vertex (3,2) to the point (1,5) you must go left 2 up 3. Find another point on the parabola by going right 2 up 3 from the vertex.

OpenStudy (zzr0ck3r):

Give the vertex we know \(y=a(x-3)^2+2\) Use the other point to solve for \(a\). \(5=a(1-3)^2+2\). Solve for \(a\) and plug that into \(y=a(x-3)^2+2\)